the normality of a solution formed by mixing 100 ml of .2 N hcl and 500 ml of .2 M h2so4 is?
 

Grade:11

3 Answers

poonam
6 years ago
Normality =molarity*n factorN=0.2*2= 0.4N1V1+N2V2= N3 (V1+V2)0.2*100+0.4*500=N3(500+100)20+200=N3*600=N3= 0.37
harsha
6 years ago
normatily=molarity*nfactor..
from the dilution law.. we get (V1+V2)=N1V1+N2V2
for hcl N1 is given and for sulphuric acid molarity is given . so basing on the structure there are 2OH groups present. So its normaility becomes 0.4.on substituing the values in the dilution law... the required answer is0.3666
harsha
6 years ago
sorry that there is a mistake in my above answer.according to the  dilution law is (V1+V2)N=N1V1+N2V2 where N1=0.2 N2=0.4 V1=100 and V2=500